A car is sold for $26,000. After one year, the value of the car is $14,300. Write an exponential function y to determine the value of the car after x years if the rate of decrease is the same each year.
y = ________
Estimate the value of the car after 4 years. Round the answer to the nearest dollar,
$ _______

Answer by stanbon(75887)

A car is sold for $26,000. After one year, the value of the car is $14,300. Write an exponential function y to determine the value of the car after x years if the rate of decrease is the same each year.
y = a*b^x
-----------
26000 = a*b^0
a = 26000
-----
14300 = 26000*b^1
------
b = 0.55
----------------
f(x) = 26000*0.55^x
----
Estimate the value of the car after 4 years. Round the answer to the nearest dollar,
f(4) = 26000*0.55^4 = $2379.16
Ans:: $2379.00
-----------------
Cheers,
Stan H.
-----------